OpenStudy (cutestgirlaround):

Tonya took out a loan to help pay for her house. She borrowed $50,000 for 15 years at a yearly simple interest rate of 5%. How much interest will she end up paying the bank? $37,500 $7,500 $375,000 $375

OpenStudy (cutestgirlaround):

@daisyduck04

OpenStudy (cutestgirlaround):

@sotoslugger

OpenStudy (r.hamza17):

Okay as you are given with the principal amount P=$50000, time period n=15 and the interest rate is r=5%, use the formula below to find the answer A=P(1+r)^n A=50000(1+0.05)^15 Now just solve it

OpenStudy (r.hamza17):

And than you an subtract the maturity amount from the principal amount to get the interset she have paid. I=A-P